A construction worker makes a 15' x 8' rectangular frame for a wall in order to check that the corners make a right angle measures the diagonal l from one corner to the opposite corner what should the diagonal measure if he built the frame properly
Answer by addingup(3677)

Pythagoras says:
a^2+b^2 = c^2
15^2+8^2 = c^2
289 = c^2 take the square root of both sides to get your answer. You should get 17.